Using LaTEX with emTEX

Leslie Lamport / Eberhard Mattes

27 October 1988 For LaTEX Version 2.09

Contents

1 Getting Started 2 1.1 Running a Sample File . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Preparing and Running LaTEX on Your Own Files . . . . . . . . . 2

2 Carrying On 3 2.1 LaTEX under OS/2 and DOS . . . . . . . . . . . . . . . . . . . . . 3 2.2 Document Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.2.1 The proc Style Option . . . . . . . . . . . . . . . . . . . . 4 2.2.2 The bezier Style Option . . . . . . . . . . . . . . . . . . 5 2.2.3 The ifthen Style Option . . . . . . . . . . . . . . . . . . 5 2.2.4 Letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.5 The showidx Style Option . . . . . . . . . . . . . . . . . . 6

2.3 Where the Files Are . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Running lablst.tex and idx.tex . . . . . . . . . . . . . . . . . 7 2.5 Differences from the Manual . . . . . . . . . . . . . . . . . . . . . 7 2.6 Using BibTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.7 Using SliTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.8 Using MakeIndex . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.9 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.10 Special Versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3 Bugs 10

4 Errata and Additions to the Manual 10 4.1 First Four Printings . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 First Five Printings . . . . . . . . . . . . . . . . . . . . . . . . . . 16

LaTEX runs on a variety of computers at many different sites. This document tells you how to use LaTEX with emTEX on computers running OS/2, MS-DOS or PC-DOS. It is not about LaTEX itself, which is described by the manual—LaTEX: A Document Preparation System, published by Addison-Wesley, available at fine book stores everywhere.

If you have a question about running LaTEX that you can’t answer by reading the manual and this document, ask Eberhard Mattes. He should also be informed of any possible LaTEX bugs or undocumented anomalies.

1 Getting Started

1.1 Running a Sample File

Before preparing your own documents, you may want to get acquainted with LaTEX by running it on a sample input file. First make your own copy of the file sample.tex by typing the following OS/2 or DOS command:

     copy c:\emtex\texinput\sample.tex

(You may have to use a different drive letter. This and all other commands are ended by typing return.) A copy of the file sample.tex is now in your current directory; you can edit it just like any other file. If you destroy or mess up your copy, typing the above command again gets you a fresh one.

Next, run LaTEX on the file sample.tex by typing:

     latex sample

When LaTEX has finished, it will have produced the file sample.dvi in your directory. You can print this file by typing a command like

     prthplj sample

See the dvidrv manual for details.

After your output has been printed, you can delete sample.dvi by typing

     del sample.dvi

1.2 Preparing and Running LaTEX on Your Own Files

You must use a text editor to prepare an input file for LaTEX. The easiest way to start learning about LaTEX is by examining the file small.tex with your text editor. You can obtain your own copy of this file, in your directory, by typing the command

     copy c:\emtex\texinput\small.tex

After you have prepared your file, whose name should have the extension tex, you must run it through LaTEX and print the output. Follow the instructions in Section 1.1, except substitute the first name of your file for “sample”. Remember to save disk space by deleting the dvi file after printing the output.

If you want to stop LaTEX in the middle of its execution, perhaps because it is printing a seemingly unending string of uninformative error messages, type Control-C (press C while holding down the key labeled CTRL). This will make LaTEX stop as if it had encountered an ordinary error, and you can return to OS/2 or DOS command level by typing X, as described in the manual.

2 Carrying On

2.1 LaTEX under OS/2 and DOS

The only special problems in using LaTEX caused by OS/2 and DOS involve the way files are handled. The first problem arises because, when a program starts to write a file, the previous version of that file is destroyed. Thus, if an error forces you to stop LaTEX prematurely (by typing Control-C ), then the files that LaTEX was writing are incomplete, and the previous complete versions have been destroyed. You probably don’t care about the output on the dvi file, but, if you are making a table of contents or using cross-referencing commands, then LaTEX also writes one or more auxiliary files that it reads the next time it processes the same input file. If the auxiliary files are incomplete because LaTEX was stopped before reaching the end of its input file, then the table of contents and cross-references will be incorrect the next time LaTEX is run on the same input file. You will have to run LaTEX a second time to get them right. If you want to avoid having to run LaTEX twice after making an error—for example, if your input is very long—then you should save copies of these auxiliary files before running LaTEX. An input file named myfile.tex and all the auxiliary files produced by LaTEX from it are included in the file specification myfile.*. Use the copy command to save copies of these files.

The second problem in using LaTEX on OS/2 and DOS involves the files that LaTEX reads. The file whose name you type with the latex command is called the root file. In addition to reading the root file, LaTEX also reads the files specified by \input and \include commands. With the OS/2 and DOS directory systems, LaTEX must know not only the names of these file but also on what directories they are. It will have no problem finding the correct files if you follow two simple rules:

1. Run LaTEX from the directory containing the root file.

2. Keep all files specified by \input and \include commands in the same directory as the root file.

If you follow these rules, you never have to type a path name when using LaTEX.

You should never break the first rule, otherwise LaTEX will have trouble finding auxiliary files. (To run LaTEX on someone else’s file, copy the file to your directory.) If you break the second rule, specifying a file from another directory in an \input or \include command, you must use a complete path name. For example, to include the file hisfile.tex from the directory \foo\bar on drive D, you can type

     \include{d:/foo/bar/hisfile}

The \ character cannot be used, so you must use / instead.

For people who don’t like to obey rules, here is exactly how LaTEX finds its files. The root file and all files specified in the LaTEX input, including documentstyle (sty) files specified by the \documentstyle command, are read from the current directory (or the specified drive/directory if the file specification includes a drive and/or a path).

LaTEX’s auxiliary files are read and written in the directory from which it is run (current directory). If LaTEX does not find a file starting in this directory, it searches all directories listed in the TEXINPUT environment variable. As a last resort, LaTEX tries to find the file in the \emtex\texinput directory of the current drive. You can change the directories in which LaTEX looks for its input files by setting the environment variable TEXINPUT. Putting the command

     set TEXINPUT=d:\mytex;d:\texinput

in your autoexec.bat or config.sys file causes LaTEX to look for files first in the current directory, then in the d:\mytex directory, then in the d:\texinput directory and finally in the \emtex\texinput directory.

2.2 Document Styles

There are six document styles and style options available that are not described in the manual: the proc style option for making camera-ready copy for conference proceedings, the bezier option for drawing curves, the ifthen option for implementing if-then-else and while-do control structures, the letter style for making letters, the showidx option for printing index entries in the margin, and the makeidx option for use with the MakeIndex program. See Section 2.8 for further information on the MakeIndex program that uses the makeidx option.

2.2.1 The proc Style Option

The proc option is used with the article document style. It produces twocolumn output for ACM and IEEE conference proceedings. The command \copyrightspace makes the blank space at the bottom of the first column of the first page, where the proceedings editor will insert a copyright notice. This command works by producing a blank footnote, so it is placed in the text of the first column. It must go after any \footnote command that generates a footnote in that column.

LaTEX automatically numbers the output pages. It’s a good idea to identify the paper on each page of output. Placing the command

     \markright{Jones---Foo}

in the preamble (before the \begin{document} command) prints “Jones—Foo” at the bottom of each page.

2.2.2 The bezier Style Option

This option defines a single command, \bezier, that draws a curved line in a picture environment. Let P_{i} be the point with coordinates (x_{i}, y_{i}), for i = 1, 2, and 3. The command

     \bezier{n}(x_{1},y_{1})(x_{2},y_{2})(x_{3},y_{3})

draws n points on the quadratic Bezier spline determined by the three points P_{1}, P_{2}, and P_{3}. The locus of points on this spline is a parabolic arc from P_{1} to P_{3} having the line P_{1}P_{2} tangent to it at P_{1} and the line P_{2}P_{3} tangent to it at P_{3}. Note that P_{2} is not on this arc unless P_{1}, P_{2}, and P_{3} are colinear, in which case the arc is a straight line. Bezier splines are useful because it’s easy to join two of them together smoothly by giving them the same tangent line where they meet.

It takes roughly 75 points per inch to form a solid line, depending upon the line thickness. See Section C.13.3 of the manual for commands to specify line thickness in a picture environment. This command is very slow, and TEX has enough memory to hold only about 1000 points plus a page of text. (Remember that TEX keeps the current page plus all as yet unprinted figures in memory.) So, the bezier command should be used for only a small number of small curves.

2.2.3 The ifthen Style Option

This option provides two programming language features that are useful only for people who already know how to program. It defines the two commands

\ifthenelse{test}{then clause}{else clause}

     \whiledo{test}{do clause}

that implement the following two Pascal language structures

if test then then clause

else else clause while test do do clause

The then, else, and do clauses are ordinary LaTEX input; test is one of the following:

• A relation between two numbers formed with <, >, or =; for example,

     \value{page}>3.

• \equal{string1 }{string2 }, which evaluates to true if string1 and string2 are the same strings of characters after all commands have been replaced by their definitions. (Upper- and lowercase letters are unequal.)

• A logical combination of the above two kinds of tests using the operators \or, \and, and \not and the parentheses \( and \)—for example:

     \not \( \value{section} = 1  \and  \equal{Jones}{\myname} \)

The test argument is a violently moving argument, which means that not only fragile commands but even some commands that are not normally fragile will break, causing TEX to enter an infinite loop. The \protect command works in these situations.

These commands, together with \renewcommand and the commands of Section C.7.4 for manipulating counters, open up a whole new world of hacking.

2.2.4 Letters

The letter document style, described in the manual, should be used for generating personal letters. Mailing labels are formatted in two columns of five 2′′ × 4-1/4′′ labels each, suitable for copying onto Avery brand, number 5352 address labels.

There are no features for making letters other than those described in the manual. However, suggestions will be accepted for such options as the ability to print ones net address as part of the letterhead.

2.2.5 The showidx Style Option

This style option, for use with the report or book document styles, causes index entries to be printed in the outer margin. It does not change the effect of \makeindex, which controls whether or not an .idx file is written. No attempt is made to avoid overprinting marginal notes. This option issues a \flushbottom declaration.

2.3 Where the Files Are

All LaTEX files mentioned in the manual, including the sty and doc files, are in the directory \emtex\texinput. Fonts are stored in two directories: the tfm files used by TEX are in \emtex\tfm, and the pixel files used by the device drivers are usually in \texfonts.

2.4 Running lablst.tex and idx.tex

A list of labels and citations in an input file is printed by running LaTEX on the input file lablst.tex, which is done by typing

     latex lablst

LaTEX will then ask for the name of the input file, which should be typed without an extension, and for the name of the main document style (e.g., article), used by that file.

The index entries on an idx file are printed by running LaTEX on the file idx.tex, which is done by typing

     latex idx

LaTEX will ask for the name of the idx file, which is typed without an extension.

2.5 Differences from the Manual

All LaTEX features described in the manual are provided by the the emTEX implementation.

2.6 Using BibTEX

BibTEX is a program for compiling a reference list for a document from a bibliographic database. It is run by typing

     bibtex myfile

where myfile.tex is the name of your LaTEX input file. This reads the file myfile.aux, which was generated when you ran LaTEX on myfile.tex, and produces the file myfile.bbl. BibTEX should be run from the directory containing myfile.tex (which should be the same directory from which LaTEX was run on that file).

If the bib file is not in the current directory or in a directory defined by the BIBINPUT environment variable or in the \emtex\bibinput directory then you must include a path as part of the file name specified by the \bibliography command. A \ character cannot appear in the argument of a \bibliography command, so you should use / instead. For example, the LaTEX command

     \bibliography{/jones/bibfiles/gnus}

specifies the file gnus.bib kept in the \jones\bibfiles directory.

There is now no formal provision for sharing bibliographic database information, nor are there programs to assist in making your own bib files. Suggestions for forming one or more common bib files are welcome.

In addition to the bibliography styles described in the manual, there is a ieeetr style that formats entries in the style of the IEEE transactions.

In addition to the usual three-letter abbreviations for the months, the following abbreviations are defined by the bibliography styles:

acmcs ACM Computing Surveys acta Acta Informatica cacm Communications of the ACM ibmjrd IBM Journal of Research and Development ibmsj IBM Systems Journal ieeese IEEE Transactions on Software Engineering ieeetc IEEE Transactions on Computers ieeetcad IEEE Transactions on Computer-Aided Design of Integrated Cir-

cuits ipl Information Processing Letters jacm Journal of the ACM jcss Journal of Computer and System Sciences scp Science of Computer Programming sicomp SIAM Journal on Computing tocs ACM Transactions on Computer Systems tods ACM Transactions on Database Systems tog ACM Transactions on Graphics toms ACM Transactions on Mathematical Software toois ACM Transactions on Office Information Systems toplas ACM Transactions on Programming Languages and Systems tcs Theoretical Computer Science

2.7 Using SliTEX

SliTEX is a version of LaTEX for making slides. To run SliTEX with a root file

myroot.tex, you type
   tex &splain myroot

(You should be connected to the directory containing myroot.tex.) Refer to Section 2.1 if you want slide files or \input files to be in a different directory from your root file.

size  default (10pt) 11pt option 12pt option
\tiny  5pt  6pt  6pt

\scriptsize  7pt  8pt  8pt
\footnotesize  8pt  9pt  10pt

\small  9pt  10pt  11pt
\normalsize  10pt  11pt  12pt
\large  12pt  12pt  14pt
\Large  14pt  14pt  17pt
\LARGE  17pt  17pt  20pt
\huge  20pt  20pt  25pt
\Huge  25pt  25pt  25pt

Table 1: Type sizes for LaTEX size-changing commands.

      \it  \bf  \sl  \sf  \sc  \tt

5pt  D  D  X  X  X  X
6pt  X  D  X  X  X  X
7pt  P  D  X  X  X  X
8pt  P  D  D  D  D  D
9pt  P  P  D  D  D  P
10pt P  P  P  P  D  P
11pt P  P  P  P  D  P
12pt P  P  P  P  D  P
14pt D  P  D  D  D  D
17pt D  P  D  D  D  D
20pt D  D  D  D  D  D
25pt X  D  X  X  X  X

Table 2: Font classes: P = preloaded, D = loaded on demand, X = unavailable.

2.8 Using MakeIndex

The MakeIndex program helps in making an index. It is described in a separate document. The MakeIndex program is actually called makeindx.

2.9 Fonts

Almost all the symbols available on our fonts can be generated by ordinary LaTEX commands. However, there are type sizes not obtainable by LaTEX’s sizechanging commands with the ordinary document styles. Consult a local TEX expert to find the TEX name for such a font.

Tables 1 and 2 allow you to determine if the font for a type style at a particular size is preloaded, loaded on demand, or unavailable. Table 1 tells you what size of type is used for each LaTEX type-size command in the various documentstyle options. For example, with the 12pt option, the \large declaration causes LaTEX to use 14pt type. Table 2 tells, for every type size, to which class of fonts each type style belongs. For example, in 14pt type, \bf uses a preloaded font and the other five type-style commands use load-on-demand fonts. Roman (\rm) and math italic (\mit) fonts are all preloaded; the \em declaration uses either italic (\it) or roman.

2.10 Special Versions

A version of LaTEX which includes German hyphenation patterns and which maps accented letters to corresponding LaTEX commands is available by typing glatex instead of latex. Accented letters do not work in the tabbing environment, use LaTEX commands instead.

3 Bugs

There are a few known bugs in LaTEX that occur very seldom and cause the user little trouble, but would be very difficult to fix. Moreover, given the nature of complex systems, it is not unlikely that the corrections would lead to even worse problems. Therefore, these bugs will probably not be fixed.

The bugs and ways to get around them are listed below. Do not worry about any of them until you are preparing the final draft, since changes to the text are very likely to cause the problem to disappear.

• In rare instances, a figure or table will be printed on the page preceding the text where the figure or table environment appears. This can be fixed by either moving the environment further towards the end of the document.

• A marginal note at the top of a page may appear in the wrong margin. This can be fixed by inserting a redundant \pagebreak command to force a page break exactly where LaTEX started the new page anyway.

• A footnote can be broken across two pages when it should fit on a single page. This happens when there is one or more figures or tables on the page. The problem is corrected by moving, towards the end of the file, the last figure or table environment that produces a figure or table on the page where the footnote starts.

4 Errata and Additions to the Manual

These are all the errors and omissions to the manual, LaTEX: A Document Preparation System reported by 31 October 1990.

4.1 First Four Printings

The following errors were corrected in the Fifth Printing. To see if your copy is one of the first four printings, turn to the copyright page—the one with the Library of Congress cataloging information, on the back of the title page. If the fifth line on that page begins “that, as of August 27, 1985”, then your copy is from the first four printings.

page xiii Add Mike Urban to the list of people thanked in the third paragraph.

page 2 In the first paragraph of Section 1.1, replace the three instances of sample.tex

by small.tex.

page 15, line 18 Replace “thay” by “that”.

page 24, line 13 Add a period to the end of the line.

page 31, line 7 Delete one “will”.

page 39, line 2 Change “six” to “five”.

page 39, line 17 Remove the extra “it”.

page 40, line −8 Change “five” to “six”.

Section 3.3.5, page 49 The name of the environment is eqnarray. There are two instances on this page of the incorrect name “eqnarry” that should be changed. page 52, last line Replace “instead of ∫ ∫ ydxdy” by “instead of ∫ ∫ zdxdy”.

pages 55–58, 124, and 173–174 Commands that define or redefine a command or environment, such as the \newcommand and \renewenvironment commands, should not be nested within one another. Doing so may result in the following TEX error:

! Illegal parameter number in definition of ... .

page 56, line −3 Change “{\bf x}” to “{x}”.

page 60, line −7 Replace “one of these environments” by “a figure or table environment”.

page 74, line 20

Change \begin{documentstyle} to \begin{document}.

page 75, line 15

Change \thebibliography to \bibliography.

page 79, line 16 Change “diplaying” to “displaying”.

page 88, line 27 Change “entire the paragraph”to “the entire paragraph”.

page 89, line 16 Change “page-breaking” to “line-breaking”.

page 95, line 5

Replace “{.01in}” by “{1.01in}”.

page 96 In the penultimate paragraph of Section 5.4.2, replace

The \vfill command is an abbreviation for \vspace{\fill}. with the following:

The \vfill command is equivalent to a blank line followed by

     \vspace{\fill}.

page 98, line 6 Remove an “i” from “directiion”.

page 102 In Figure 5.1, replace the two occurrences of “-1.8” by “−1.8”.

page 104, line −3 Correct the spelling of “positioning”.

page 113 Change the label on the arrow above Item 2 from \itemsep to

\itemsep + \parsep.

page 116, line 20 Change “fourteen-point Plus Roman” to “twelve-point Plus Roman”.

page 118, line −10 Change this line to:

... (myfile.tex [1] [2] [3] (part1.tex [4] [5]) (part2.tex [6] [7]

page 129 The following LaTEX warning messages should be listed:

     No ...  typeface in this size, using ...

A type style declaration specified a type style and size combination that is not available, so LaTEX is substituting another one.

     ...  in math mode.

The indicated command is not permitted in math mode but was used there. Remember that \boldmath, \unboldmath, and size-changing commands may not be used in math mode. pages 141–142 Change “von Beethoven” to “van Beethoven”.

page 142 Add the following near the bottom of the page, just above the Titles heading.

If an author or editor field has more names than you want to type, just end the list of names with and others; the standard styles convert this to the conventional et al.

page 147, line 6 Change “Scribe-compatible” to “Scribe-compatible”.

page 152, line −6 Add \newtheorem and \hyphenation to the list of global declarations.

page 157, line −10 The first paragraph of Section C.3.1 should note that the *-form of a sectioning command does not affect the running head.

page 158, line −12 Change “chapters” to “appendix chapters”.

page 160, top line Change “(Section 5.6)” to “(Section 5.3)”.

page 160, line −9 Change the description of the openbib style option to: “Causes the bibliography (Section 4.3) to be formatted in open style. (See van Leunen [7].)”

page 163, line 19 Change “Section 5.1.2” to “Section 5.1.1”.

page 168, line −5 Replace “printed as \␣” by “printed as ␣”. page 169, line −5 Change “first and third rows” to “first and third columns”.

page 169, line −3 Add a period to the end of the sentence.

page 170 Add the following paragraph after line 5:

An overfull \hbox warning occurs if a formula extends beyond the prevailing margins. However, if the formula does lie within the margins, no warning is generated even if it extends far enough to overprint the equation number.

page 174, line −1 Add the following sentence: “It is a global declaration.”

page 176 Change the four lines immediately following the heading for Section C.8.1 to:

      \begin{figure}[loc] body \end{figure}
      \begin{figure*}[loc] body \end{figure*}
      \begin{table}[loc]  body \end{table}
      \begin{table*}[loc] body \end{table*}

page 185, line −5 Replace “suppresses command” by “command suppresses”.

page 187, lines 15–17 The sentence “It also writes bib files… ” is redundant and can be eliminated.

page 188, lines 1, 2, 4, 8 Change “key lst” to “key list”.

page 191, line 13 The word “paragraph” misspelled. page 191, line 14 Remove the space between \begin and {sloppypar}. (This is for consistency only; LaTEX ignores the space.)

page 199 Add the following sentence to the last paragraph on the page:

Words typeset in typewriter style or in two different styles are not hyphenated except where permitted by \- commands.

(This is a change to LaTEX made on 18 December 1985.) Also, add the following index entry citations to this page: “\-”, “hyphenation, suppressed”, and “typewriter type style, no hyphenation in”.

page 217 Add the subentry “openbib, 160” to the index entry “document-style option”.

page 223, column 2, line 13 Change “mad” to “made”.

page 225 Add page 47 to the index entry for “Leunen, Mary-Claire van”.

page 228 In index entry for \multicolumn, change “194” to “184”.

Tear-Out Command Sheet In the first column, sixth line after “Sentences and Paragraphs” heading, replace “& &” by “& \&”.

4.2 First Five Printings

page xiii, line −7 The hyphenation of “Leunen” is obviously incorrect.

page7, line 33 Add the following sentence after “[6].”

(Many of the details of LaTEX, including command names, were also taken from Scribe.) page 12, line −7

Replace “every” with “ever”.

page 14

The discussion of space after a period should include an example like et␣``al.''\␣ showing where to put the ␣ command when there is no space right after the period.

Section 3.3.3, page 47

Somewhere, it should be mentioned that each item in an array environment is a separate math formula.

page 50, line 5

Replace the sentence beginning “A + or −” with:

A + or − that begins a formula (or certain subformulas) is assumed to be a unary operator, so typing $-x$ produces −x (and typing $\sum -x_{i}$ produces ∑ −x_{i}), with no space between the “−” and the “x”.

page 53, line 16

Change “math itallic is the default in math mode” to “math italic is almost identical to the default type style of math mode”.

page 71, last line

Change “comand” to “command”.

page 84

Add the following paragraph just after line 2.

Some commands, such as \chapter, change the style of the current page. You can change it yourself with the \thispagestyle command.

Also, add the appropriate index entries for \thispagestyle and \chapter.

page 88, line 29

Change “its” to “it’s”. page 91, line 5 It should be noted that in the book document style, \chapter uses

\cleardoublepage.

page 111, line −4 Remove the extra t from “withtout”.

page 114, line 31 Change “Section 5.6” to “Section 5.3”.

page 122, line 17 Change the sentence that begins “The most probable cause is” to the following:

The most probable cause is a list-making environment with no \item command.

Page 126, line 12ff An un\protected command—even a robust one—in the argument of one of these commands can cause TEX to run out of this kind of space. It’s a good idea to \protect all commands in these arguments.

Page 153, top of page It should be noted that if the invisible “word” occurs at the end of a paragraph, not attached to a real word, then it could appear on a line by itself, producing a blank line in the output.

Page 156 In the third line of the description of \footnotemark, remove the phrase “just

like \footnote”.

Page 160, line −7

Replace “equations” with “equation”.

Page 162 In the description of \pagenumbering, it should be mentioned that the command sets the value of the page counter to 1.

Page 161, line −9 Replace “head information” by “head information (but not page number)”

Page 162, line 6

Replace “heading” by “headings”.

Page 173, line −2 Replace “cmd” with “\begin{nam}”.

Page 177, line 1 It should be noted that, in a two-column page style, a single-column figure can come before an earlier double-column figure, and vice-versa.

Page 177, line 3 Replace “pos” by “loc”.

Page 177, line −1 Replace “page” with “text page”.

Page 178, lines 1 and 8 Replace “page” with “text page”.

Page 181, line 11 Replace “or \ command” with “or \' command”.

page 183 The description of \extracolsep should mention that no extra space is put to the left of the first column.

page 191 Add to the description of \pagebreak the information that it has an effect only when used in paragraph mode, not inside a parbox.

page 191, line −2 Add a “t” to “explicily”. page 213 Change the page reference for “\bibitem, moving argument of” from 153 to 152.

page 211 Change index entry for \AE to page 40.

page 213 Change the first page reference for “bbl (bibliography) file” from 74 to 75.

Add page 95 to the entry for “backspace”.

page 217 Add an index entry “diacritical marks, see accents”.

Move the index entry for \def to its proper place on the preceding page.

page 219, index entry for “figures” Remove the “and” from the subentry for “output by \end{document}”.

page 219, column 1, line 3 Replace “anomolous” by “anomalous”.

page 223, left column, line 2 Remove the page number “79” from the first subentry under \index.

page 228 Change the page reference for “\multicolumn” from 194 to 184.

page 228 Remove the extra “,” from the index entries for \o and \O.

page 234 Add “48–49” to the index entry for \right.